Extracted Ticket


Ticket Number  : SUNETTICKET-6453
Ticket Type    : Unscheduled
Ticket Status  : Closed
Ticket Summary : Link Degradation STHB-KOB1
Ticket Scope   : Link

Ticket Opened  : 20200625 14:10 UTC
Ticket Closed  : 20200730 5:50 UTC

Problem Start  : 20200625 11:41 UTC
Problem End    : 20200729 16:48 UTC

Affected organisations:

      * NORDUnet, SUNET, Netnod
        ----

Description:

      * We see higher attenuation between the fiber in Stockholm
        (STHB) and Kolbotten (KOB1)
        ----

Impact:

      * Reduced capacity for affected organizations. 
        SUNET core link between Nyköping and Stockholm is
        affected.
        
        Below services to our customers are affected as well.
        
        SU-S003624	DWDM - Alienwave	STHB - ORE NDN2
        193.60THz	NORDUnet 
        SU-S003638	DWDM - Alienwave	STHB- ORE NDN3 193.20THz
        (Cenia)	NORDUnet 
        SU-S003649	DWDM - Alienwave	STHB - ORE NDN2
        192.55THz	NORDUnet 
        SU-S003656	DWDM - Alienwave	Wave for Netnod
        "dwdm1-stb.sth.netnod.se" to
        "dwdm1-nnb.mmo.netnod.se"	Netnod
        SU-S003659	DWDM - Alienwave	STHB - ORE NDN1
        193.10THz	NORDUnet
        ----

Update:

      * 20200625 14:22 UTC
        We are investigating the issue.
        ----
        
      * 20200625 14:29 UTC
        We are going to clean the fiber in Stockholm node to see if
        the values get better. The services will go down during the
        troubleshooting.
        ----
        
      * 20200625 14:36 UTC
        The services have restored after the cleaning. It did not
        help. SUNET will escalate it to its fiber provider.
        ----
        
      * 20200625 15:13 UTC
        The issue is escalated to our fiber provider.
        ----
        
      * 20200630 12:15 UTC
        Asked fiber provider for an update.
        ----
        
      * 20200720 10:41 UTC
        A maintenance date will be fixed by the fiber provider for
        checking the fiber.
        ----
        
      * 20200729 17:25 UTC
        The issue with the fiber has been fixed and we are observing
        better signal.
        ----
        
        
Final ticket report:

      * We experienced higher attenuation between the fiber between
        Stockholm (STHB) and Kolbotten (KOB1). This was due to a bad
        patch at an ODF. The patch is now replaced and the
        attenuation is fixed.
        ----